The demand and supply curves for the market are given by: Demand: Qd = 16000 – 24P Supply: Qs = 2000 + 32P

A maximum price of $200 is proposed.

(ii) Calculate the reservation price for the seller (to the nearest whole number)
Answer #1

The supply equation is given as 2000+32P

The reservation price would be

Set QS=0,

0 = 2000+32P

32P = 2000-0

P = 2000/32 = 62.5

The reservation price for the seller = 62.5